Equivalent Units of Materials Cost

The Filling Department of Eve Cosmetics Company had 2,100 ounces in beginning work in process inventory (50% complete). During the period, 34,700 ounces were completed. The ending work in process inventory was 1,700 ounces (20% complete).

What are the total equivalent units for direct materials if materials are added at the beginning of the process?
units

PART 2

Equivalent Units of Conversion Costs

The Filling Department of Eve Cosmetics Company had 6,100 ounces in beginning work in process inventory (20% complete). During the period, 67,600 ounces were completed. The ending work in process inventory was 6,100 ounces (60% complete).

What are the total equivalent units for conversion costs?

If required, round to the nearest unit.

units
Answer #1

Equivalent unit of production = Unit completed+(Ending work in process*percent completion)

1) Equivalent unit of material = 34700+(1700*100%) = 36400 Units

2) Equivalent unit of conversion = 67600+(6100*60%) = 71260 Units